OpenStudy (anonymous):

Let * be a binary operation on Real numbers defined by a*b=ab/4 Given a*b=9 Find a and b

OpenStudy (anonymous):

help ! @TuringTest , @Mani_Jha , @dumbcow ,

mathslover (mathslover):

ok so we are given with the following information : a * b = ab/4 a*b = 9 this means \(\large{\frac{ab}{4} = 9}\) , right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

---> ab =36

mathslover (mathslover):

Now , can you find what will be the value of "ab" ?

OpenStudy (anonymous):

ab=36 ?

OpenStudy (anonymous):

100010 =36

OpenStudy (anonymous):

but how i find value of a and b ?

mathslover (mathslover):

Yes right @aman07

mathslover (mathslover):

OK let me think for what to do after this, wait!

OpenStudy (anonymous):

ok

mathslover (mathslover):

What we have yet as a first way to get rid of this is "hit and trial" but really I don't think that hit and trial will be a satisfactory method ....

OpenStudy (anonymous):

well i knw to do that ..but there is some method using identity element and stuff

mathslover (mathslover):

http://www.math.csusb.edu/notes/binop/node2.html I saw something here

OpenStudy (anonymous):

:( i have that already in text.. i am stuck in this particular qn :(

OpenStudy (anonymous):

i got it now. you have ab =36 that means a=3 and b =6. in binary system, 3= 10 (one zero) and 6 = 1001. make multiplication between them , you got exactly 36 in both code (binary and 10 system)

OpenStudy (anonymous):

wth ??

OpenStudy (goformit100):

ok

mathslover (mathslover):

I tried to search some relative problems on binary operations on internet and I found this question : if : \(\large{a\alpha b = |a-b| }\) then what will be \(\large{6\alpha 8 }\) . Of course that will be 2 but the question given is not exactly what we have in m

mathslover (mathslover):

In my question (contd. from the previous post.) Ok, I have never learnt it but I thought it is a type of mental ability question... Is it also related to binary numbers ?

OpenStudy (anonymous):